OpenStudy (anonymous):

Will give a medal to the best answer please show work :) a+s=150 $45.00a+$28.50s=$4770

OpenStudy (agent0smith):

You could use substitution or elimination... do you prefer one or the other?

OpenStudy (anonymous):

Doesn't matter :)

OpenStudy (anonymous):

From the first equation, a = 150 - s (subtracting 150 from both sides) Now, substitute a = 150 - s into the second equation, so you have 45(150 - s) + 28.50 s = 4770 6750 - 45s + 28.50 s = 4770 6750 - 16.5s = 4770 -16.5s = -1980 s = 120. We know s = 120; to find a, substitute s = 120 into the first equation, a + s = 150 120 + s = 150 s = 30 Therefore, the final answer is a = 120 and s = 30.

OpenStudy (agent0smith):

k then, elimination is easy enough... multiply the first equation by 45: 45a+45s=45*150 $45.00a+$28.50s=$4770 now subtract the 2nd equation from the first.

OpenStudy (anonymous):

Thanks :) I wish I could give more than one medal but I love both your answers I like yours better though agent :p